11.3.35 Q10 A pet store has 13 ​puppies, including 3 ​poodles, 5 ​terriers, and 5 retrievers. If Rebecka and​ Aaron, in that​ order, each select one puppy at random with replacement​ (they may both select the same​ one), find the probability that Rebecka selects a terrier and Aaron selects a retriever. The probability is____
Answer #1

Rebecka select a terrier = 5/13

Aaron selects a retriever = 5/13

Probability for both event = 5/13 * 5/13 = 25/169

= 0.148

Please revert back in case of any doubt.

Please upvote. Thanks in advance.